OpenStudy (anonymous):

how to do this? 3x-4y=-1 and x+y=5 in substitution

OpenStudy (anonymous):

haha. okay, so apparently the first answer wasn't good. you need to solve one of your equations for one of the variables. for example, solving the second equation for "y" gives y = 5 - x

OpenStudy (anonymous):

this can be substituted. to make this easier, put the right hand side of your solved equation directly into the left, using parentheses, like this: 3x - 4(5-x)=-1

OpenStudy (anonymous):

okey.okey okey yes i forgot .tnx

OpenStudy (anonymous):

y = 5 -x 3x - 4y = -1 3x - 4(5-x) = -1 (see azeotrope's explanation) 3x -20 + 4x = -1 distribute -4 to the terms inside the parenthesis 7x -20 = -1 combine similar terms 7x = -1 + 20 transpose -20 to the other side 7x = 19 do the operation x = 19/7 divide both sides by 7